Circular No. 7312

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE J. Tonry, University of Hawaii; N. B. Suntzeff, Cerro Tololo Inter-

american Observatory (CTIO); and S. Jha, Harvard-Smithsonian Center for Astrophysics (CfA), on behalf of the High-Z Supernova Search Team (cf. IAUC 6160, 6646), report the discovery of 20 new supernovae on CCD images taken with the 3.6-m Canada-France-Hawaii Telescope (CFHT; + 12k mosaic camera) on Nov. 2 and 3 by P. M. Garnavich, B. Barris, B. P. Schmidt, A. M. Soderberg, Tonry, and Jha, by subtracting CFHT images of the same fields taken on Oct. 3, 4, 7, and 8 by Garnavich and G. Wilson. The supernovae were confirmed in CCD images taken with the CTIO 4-m Blanco Telescope (+ Mosaic II imager) on Nov. 6 and 7 by R. A. Schommer, R. C. Smith, L. G. Strolger, A. Clocchiatti, P. Challis, Suntzeff, and Jha, and were further confirmed with spectra taken with the Keck-2 telescope (+ LRIS) on Nov. 8, 9, 12, 13, and 14 by A. V. Filippenko, A. Coil, A. G. Riess, R. P. Kirshner, Tonry, Garnavich, and Jha. Each supernova is within 2′′ of the host galaxy (no clear host is seen for SN 1999fv).

SN  1999 UT  α_{2000}  δ_{2000}  I  z Type
1999ff Nov. 2 2^{h}33^{m}54.^{s}39 + 0^{o}32′55′′.6 23.1 0.46 Ia
1999fg Nov. 3 2 27 41.88 + 0 39 13.3 22.9 0.42 ?
1999fh Nov. 3 2 27 58.33 + 0 39 36.8 22.4 0.36 Ia
1999fi Nov. 3 2 28 11.67 + 0 43 39.3 24.1 0.78 Ia
1999fj Nov. 3 2 28 23.72 + 0 39 09.6 23.5 0.81 Ia
1999fk Nov. 3 2 28 53.88 + 1 16 24.2 24.3 1.05 Ia
1999fl Nov. 3 2 30 05.53 + 0 44 52.6 24.9 0.3 II
1999fm Nov. 3 2 30 35.62 + 1 09 43.3 23.6 0.93 Ia
1999fn Nov. 3 4 14 03.88 + 4 17 55.0 23.5 0.47 Ia
1999fo Nov. 3 4 14 45.75 + 6 38 34.3 24.0 1.07 Ia
1999fp Nov. 3 4 15 02.59 + 4 21 46.4 24.1 0.34 II?
1999fq Nov. 3 23 25 46.50 + 0 17 49.2 21.7 ?  ?
1999fr Nov. 3 23 27 16.38 + 0 03 46.8 23.7 ?  II?
1999fs Nov. 3 23 27 17.17 + 0 07 50.3 21.3 ?  ?
1999ft Nov. 3 23 28 14.90 + 0 05 06.6 24.7 0.69 Ia
1999fu Nov. 3 23 29 48.21 + 0 08 27.0 24.9 1.03 Ia
1999fv Nov. 3 23 30 35.80 + 0 16 40.0 24.5 1.23 Ia
1999fw Nov. 3 23 31 53.03 + 0 09 32.3 21.1 0.28 Ia
1999fx Nov. 3 23 33 37.79 + 0 12 03.8 23.1 0.49 Ia?
1999fy Nov. 3 23 33 46.90 + 0 12 16.0 22.7 0.28 II?

1999 November 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green